Don't copy from the other answer! Please write the result for part 2

  1. Substitution effect

The Slutsky equation decomposes a change in consumption caused by a price change (income effect and substitution effect). Find the substitution effect of a price change in the following cases:

  1. εU= -0.7, εy =1.4 and budget share (b) = 0.2

  1. εU= -0.9, εy = 0.8 and U = x10.5x20.5
Answer #1

a) the intire change in utility will be because sub situation effect. ep=ep^h-e , ep=ep^h-eyb where ep=uncompensated price elasticity, ep^h=compensated price elasticity, ey=income elasticity, b=budget share ep^h=(-0.7)+1.4x0. 2=-0.42., ep^h=-0.42 thank you